### v1

# Stronger Than You Think: Benchmarking Weak Supervision on Realistic Tasks

Abstract
Weak supervision (WS) is a popular approach for label-efficient learning, leveraging diverse sources of noisy but inexpensive weak labels to automatically annotate training data. Despite its wide usage, WS and its practical value are challenging to benchmark due to the many knobs in its setup, including: data sources, labeling functions (LFs), aggregation techniques (called label models), and end model pipelines. Existing evaluation suites tend to be limited, focusing on particular components or specialized use cases. Moreover, they often involve simplistic benchmark tasks or de-facto LF sets that are suboptimally written, producing insights that may not generalize to real-world settings. We address these limitations by introducing a new benchmark, BoxWRENCH, [^1] The box wrench is the most ubiquitous and practical wrench. designed to more accurately reflect real-world usages of WS. This benchmark features tasks with (1) higher class cardinality and imbalance, (2) notable domain expertise requirements, and (3) multilingual variations across parallel corpora. For all tasks, LFs are written using a careful procedure aimed at mimicking real-world settings. In contrast to existing WS benchmarks, we show that supervised learning requires substantial amounts (1000+) of labeled examples to match WS in many settings.
1 Introduction
Weak supervision (WS) aims to address the labeled data bottleneck for supervised machine learning. It uses multiple weak but inexpensive sources of signal and combines them into high-quality pseudolabels that can be used for training downstream models [32, 33, 37]. These weak sources can be diverse, including but not limited to: heuristic rules encoded into small programs, queries to knowledge bases, and pretrained models. Frameworks implementing WS are hugely popular and are widely applied in industry [3, 34] and academic settings [12, 40].
WS frameworks typically have a simple three-stage approach. First, they formalize weak sources into labeling functions (LFs). In contrast to manual labeling, these can be automatically applied to an entire unlabeled dataset. Next, since LFs are inherently noisy and may conflict with one another, a label model (LM) is used to estimate the quality of each source (typically without access ground truth labels) and then to aggregate their outputs into high-quality pseudolabels. Finally, these pseudolabels can be used to train a downstream model. A vast literature studies variations on this basic recipe, with diverse approaches to crafting LFs, creating LMs, and noise-aware training of end-models [45].
For practitioners, a key question is when is WS useful? While it is natural to produce benchmarks that answer this, surprisingly, there has been relatively little work doing so. One reason for this may be the overall complexity of WS pipelines. The performance of a WS system varies with (1) the underlying task and data, (2) the LFs, (3) the choice of LM, and (4) the choice of end model and training procedure. Several benchmarks predominantly focus on only one of these. For example, WRENCH [44] focuses on evaluating (3), the LM, while AutoWS-Bench-101 [35] focuses primarily on (2), the LFs, and specifically, techniques for automatically generating model-based LFs.
Recently, Zhu et al. [47] tackles the goal of quantifying the value of WS. They argue that the benefits of WS are often overestimated by showing that fine-tuning on only 50 ground-truth labels can achieve comparable—or better—results than certain WS approaches for many benchmark datasets. They suggest that WS may not be broadly useful, as obtaining 50 “clean” labels is rarely prohibitive, and data at this scale (or larger) may still be needed for tuning or evaluation even when using WS.
In this work, we show that these findings result from the simplicity of existing datasets rather than the inherent weakness of WS. In particular, we identify two main issues with the current WS benchmarks that led to this result, and show that WS may be stronger than is thought in more realistic settings:
  1. 1.
    Benchmark datasets usually have too few classes, are balanced, or aren’t specialized enough to be representative of real-world datasets.
  2. 2.
    WS depends on the quality of LFs, and LFs from current benchmarks can be improved.
We introduce a new benchmark, BoxWRENCH, that addresses these two challenges. It enables us to quantify the practical advantages of WS in a wide range of settings. Our findings indicate that even simple WS approaches often provide substantial value. We address the issues we identified by
  • Proposing new WS benchmarks based upon tasks that involve high-cardinality label spaces, imbalanced classes, and/or require specific domain knowledge.
  • Showing that by adhering to careful LF design practices, we can write effective LFs for these tasks that can even improve upon existing benchmark LFs.
Additionally, we experiment with the re-usability of LFs across related task specifications, another key practical advantage of WS deployments that does not show up in standard benchmarks. Using the MASSIVE dataset [10], we introduce a new method to reuse existing LFs on multi-language datasets, providing an out-of-the-box performance boost.
Our benchmark consists of five text-classification WS tasks that showcase the power of WS in a variety of challenging real-world scenarios. For two of our tasks, we produce new LFs, while for one, we improve the existing LFs from WRENCH [44]. The design of these LFs follows a rigorous procedure that we release as part of our benchmark, acting as guidance for LF design and for WS benchmarking overall. We publicly release all experiment code and datasets. [^2] https://github.com/jeffreywpli/stronger-than-you-think
2 Background and Related Work
We provide a brief background on WS techniques and benchmarking efforts. We note in passing that the term WS can be overloaded and is also applied to other families of techniques, particularly in computer vision [27, 20].
Here, we use WS to refer to approaches that fall under programmatic WS [32, 45].
Weak Supervision. WS uses imperfect label sources, often small programs, as LFs to generate labels for unlabeled data. Perhaps the most popular types of LFs are heuristics or rules obtained from subject matter experts [31] encoded into programs. Other potential sources include knowledge base queries, pretrained models, and more [16, 21, 8, 28, 17]. Many works focus on either improving the aggregation technique (the LM) or obtaining improved approaches to crafting LFs. Other methods aim to improve the end model learning algorithms when training on WS pseudo-labels (e.g. COSINE [43]). The simplest aggregation technique is directly using majority vote. Others seek to learn (without ground truth labels) the accuracy of the LFs and thus perform a higher-quality aggregation [31, 13, 38]. For LF construction, variations such as learning small models on tiny amounts of data [41], or using code-generating large language models to craft LFs have been proposed [19, 18, 15].
WS Benchmarks.
Existing WS benchmarks typically focus on a particular component of a WS system or a particular use case. WRENCH [44] benchmarks LMs and is therefore aimed at aggregation. AutoWS-Bench-101 [35], in contrast, studies the effectiveness of automated LF construction techniques. Finally, WALNUT [46] studies WS techniques in the context of natural language understanding. All of these benchmarks are highly useful, but do not attempt to measure the value of WS techniques more broadly. A recent effort by Zhu et al. [47] tackles this question and finds that in particular settings, WS may not be of great value. Specifically, it suggests that only a small amount of labeled data is sufficient to train a supervised model to a level of quality equivalent to that provided by WS. We are inspired by this work, studying whether we can obtain similar findings across a broader range of realistic scenarios.
3 Methodology and Datasets
We establish the goals, problem setting, datasets, and experimental setup used in BoxWRENCH.
3.1 Goals
The ultimate goal of BoxWRENCH is to bridge WS research and practice by introducing more realistic benchmarks for WS. The first step towards such a goal is to gain a better understanding of the question: when is WS useful? To do so, we first gather a suite of datasets which addresses two key areas in which current WS benchmarks fall short.
  1. 1.
    Benchmark datasets tend be simplistic, exhibiting properties that are not representative of many real-world problems. This includes having a small label space (often binary), balanced label distributions, and relying on general rather than domain-specific knowledge.
  2. 2.
    Current WS benchmarks are used with de-facto LF sets that vary in quality. A poorly written LF set may also result in a less realistic benchmark (e.g., if a task involves domain expertise but experts were not involved in writing the LFs).
To address the first issue, we introduce WS tasks that directly target the aforementioned gaps: focusing on those with greater class counts, class imbalance, and domain-specificity. To address the second issue, we place care into writing higher-quality LFs for all datasets, including improving existing LFs. Using these datasets, we aim to measure how many labeled examples are needed before supervised learning catches up to WS techniques. We formalize this notion by plotting the performance curves of WS techniques and fully supervised learning (as functions of the number of labels) and analyzing the corssover points where these curves intersect. To show that WS is effective, the crossover point in which supervised learning surpasses WS should be sufficiently high, i.e., WS cannot be matched by simply labeling a trivial amount of examples. Using crossover points to measure the effectiveness of WS, we aim to establish a regime in which WS is practically useful on our suite of more challenging and realistic datasets.
3.2 Problem Formulation
Let $\mathcal{D}\subseteq\mathcal{X}\times\mathcal{Y}$ be our data distribution. We first sample an unlabeled training set with $n$ examples $X_{\text{train}}=[x_{i}]_{i=1}^{n}$, and a small labeled validation set of $m$ examples: $D_{\text{val}}=[X_{\text{val}},Y_{\text{val}}]$ with $X_{\text{val}}=[x_{i}]_{i=n+1}^{n+m}$ and $Y_{\text{val}}=[y_{i}]_{i=n+1}^{n+m}$ with $x_{i}\in\mathcal{X}$ and $y_{i}\in\mathcal{Y}$. We are interested in learning a function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ that minimizes the expected risk $R(f_{\theta})=\mathbb{E}_{(x,y)\sim\mathcal{D}}[L(f_{\theta}(x),y)]$ where $L$ is a loss function and $\theta$ are the model parameters. We then evaluate $f_{\theta}$ on an i.i.d. drawn test set, $D_{\text{test}}$. In the case of WS, let $\mathcal{H}$ be the hypothesis class of the LFs. For every $h\in\mathcal{H}$, we have $h:\mathcal{X}\rightarrow\mathcal{Y}\cup\{-1\}$, where $-1$ denotes an abstention. Let $\text{LF}_{\text{ws}}\subset\mathcal{H}$ be the set of LFs that we use to generate LF votes. For every $x_{i}\in X_{\text{train}}$, $\lambda_{i,j}=h_{j}(x_{i})$, where $h_{j}\in\text{LF}_{\text{ws}}$, is $j$-th LF votes for $x_{i}$. Using an LM that aggregates the LF votes, we obtain $\widehat{y}_{i}=\text{LM}\left([\lambda_{i,j}]_{j\in[|\text{LF}_{\text{WS}}|]}\right)$, namely, the weak label for $x_{i}$. We construct the weak labels for the training set as $\widehat{Y}_{\text{train}}=[\widehat{y}_{i}]_{i=1}^{n}$. We then use $D_{\text{WS}}=[X_{\text{train}},\widehat{Y}_{\text{train}}]$ as the weakly labeled training set, to train a model $f_{\theta_{\text{WS}}}$.
Following the continuous fine-tuning (CFT) setting of Zhu et al. [47], we start with a pre-trained model $f_{\theta_{\text{WS}}}$ and fine-tune it on the clean validation set $D_{\text{val}}$ to obtain $f_{\theta_{\text{CFT}}}$. An important baseline is a supervised-only approach, where we train the model directly on labels from the clean validation set $D_{\text{val}}$ to obtain $f_{\theta_{\text{SUP}}}$.
3.3 Datasets
Existing WS benchmarks often exhibit low class cardinality (under 5) or use label distributions that are highly balanced (i.e. near uniform). However, real-world WS tasks often do not conform to these assumptions, as they often involve imbalanced label distributions and high class cardinality. To evaluate WS in realistic regimes, we propose a collection of datasets [^3] With the current exception of Amazon31, which was recently retracted, all of the datasets we use are publicly available and with proper licenses (see Appendix B) with varying levels of difficulty and requirements for domain expertise, having a range of class distributions and cardinalities. We first describe each of the datasets used in BoxWRENCH, with their metadata shown in Table 1, and describe our LF design procedure.
Table 1: The datasets used in BoxWRENCH  and their metadata. For MASSIVE, the dataset sizes are the amounts per available language.
Dataset Class Train Valid Test
Banking77 77 9,003 1,000 3,080
ChemProt 10 12,600 1,607 1,607
Claude9 9 5,469 200 2057
[…]
Amazon31 31 131,781 5,805 17,402
  • Banking77 [5, 23] comprises online banking queries annotated with one of 77 user intents. It has 209 keyword-based LFs.
  • ChemProt [22] is a chemical relation classification dataset comprising 1,820 PubMed abstracts with chemical-protein interactions annotated by domain experts. The dataset was studied in [2, 47]. Previous works on WS created LFs for the dataset and showed the efficacy of WS in the dataset [43]. We push the boundaries of WS in the dataset by modifying the LFs to incorporate the distance between the chemical and protein entities in the text, among other minor modifications (see Fig. 13).
  • Claude9444Claude9 is imbalanced, with 90%+ of data belonging to one class, so we evaluate using macro-averaged F1. is based on UNFAIR-ToS [7, 25], which includes 50 Terms of Service (ToS) from online platforms and sentence-level annotations with 8 types of unfair contractual terms (potentially violating user rights according to EU consumer law). [^5] Art.3 of Direct. 93/13, Unfair Terms in Consumer Contracts (http://data.europa.eu/eli/dir/1993/13/oj). LFs for this dataset were created by a law graduate student.
  • MASSIVE18 [10, 23] is a parallel corpus of 52 languages with annotations for Natural Language Understanding tasks of intent prediction and slot annotation. The dataset is parallel, because given any two languages subsets, there exist a bijection between the examples in the two subsets. We construct several WS datasets based on the original MASSIVE, we also provide pipeline that can generate similar datasets in different languages easily.
    • MASSIVE60 [10, 23] we created a high cardinality version of MASSIVE18 that has weak labels by splitting existing classes of MASSIVE18 into detailed sub classes. The original LFs have been generalized, with a mechanism that selects a random subset from the original results for each instance (see Fig. 9).
  • Amazon31 is built from the Amazon product reviews [1] dataset consisting of reviews and their categories. [^6] This dataset has been taken down. We include it in experiments but we will not release it in BoxWRENCH. Due to each class’s high overlap and conflict rate, we merged several labels and reduced the class cardinality to 31.
LF Design Pipeline.
We randomly selected samples with clean labels from the training set to create a development set for LFs. We use 250 examples as a development set for Amazon31, which is consistent with the development of LFs for Banking77 and MASSIVE18 in [23]. For Claude9, the development set had 24 examples. We manually inspected labeled examples in the development set, and identified patterns for each class. Then we created multiple keyword, dictionary-based, and regular expression-based LFs for each class. We calculated LF statistics (which do not require label information) on the original training and validation sets, including coverage and LF conflict ratios. To to evaluate the final LFs, we calculate their accuracy scores on the original validation set. [^7] We created the LFs ourselves and did not hire annotators.
Figure 1: We compare end models trained with three different pipelines: (1) Supervised, which fine-tunesuses clean labels from the validation set for fine-tuning; (2) Weakly Supervised, which applies all LFs and aggregates them with the LM then fine-tuning an EM using the training data with the aggregated labels; using data with clean labels from the validation set for hyperparameter searching or early stopping. Continuous-Fine-Tuning: using parameters of the EMs after WS experiments, and then continuously fine-tuning them on the same data with clean validation labels.
Figure 2: Using an off-the-shelf translator (DeepL) to reuse English LFs for the multilingual variants of MASSIVE18. Our approach is to apply the existing LFs written for MASSIVE18-En by first translating non-English versions of the dataset to English.
3.4 Experimental Setup
We describe various aspects of our experimental setup: our pipeline, learning scenarios, a description of how we evaluate when WS is useful.
Pipeline. For backward compatibility, our experimental pipeline is based on WRENCH  [44]. This allows users to experiment with different LMs, EMs, and LFs in a standardized format, and with WS-specific tooling for dataset manipulation. We use the majority vote (MV) as the default LM for all of our WS experiments (following WRENCH’s takeaway, MV produce good if not the best performance on many of the benchmarking datasets). Additionally, we test the more complex COSINE [43] and ARS2 [39], which did not consistently improve over MV (see Appendix I).
For the end-model, we fine-tune a pretrained RoBERTa model in most cases [26] and also try BERT variants for domain-specific (i.e., LegalBERT[6] for Claude9 and Sci-BERT[4] for ChemProt) and non-English tasks (BERT-base-chinese, NB-BERT-base, and BERT-base-japanese [14, 29, 30] for MASSIVE18). Following the procedure used by Zhu et al. [47], we randomly select a set of hyperparameters from their provided hyperparameter search spaces. We finetune for a fixed 6,000 total steps and average each experiment over 5 trials. Notably, we can successfully replicate the results of [47] with our setup on existing WS benchmarks, as shown in Appendix E.
We run the experiments on NVIDIA A100, A40, A6000, A4000, and RTX-4090 GPUs.
Learning Scenarios. We include the three types of learning scenarios that we compare: supervised learning, WS learning, and CFT, all involving clean labels coming from only the validation set (as shown in Fig. 1).
  • Supervised: We use clean labels from the validation set directly for fine-tuning.
  • Weakly Supervised: We aggregate all weak labels from the training data with an LM, then fine-tune an EM using the training data with the aggregated labels. A standard WS pipeline uses data with clean labels from the validation set for hyperparameter searching or early stopping. This experiment varies the size of validation data for early stopping.
  • Continuous-Fine-Tuning: We further fine-tune the models trained using only weak supervision on the same set of clean validation labels.
Crossover Points. We solidify the concept of WS being useful by studying the ‘crossover points’ in performance over the validation set sizes, when comparing WS to other techniques. In our setting, crossover points we care about are between Supervised method and the WS/CFT. Crossover points, as an object of study, underpin our experimental results comparing WS to supervised learning.
4 Results and Analysis
In this section, we present our main results and analysis. In Section 4.1, we first investigate the crossover points for our new datasets and compare them with existing ones. In Section 4.2, we show that crossover points can be significantly increased when LFs are written more carefully. Section 4.3 showcases another dimension of the usefulness of WS, as we demonstrate how LFs can be adapted in a multilingual setting. Finally, we study whether different LMs perform better on our more challenging new tasks in 4.4.
4.1 Crossover Points
We define the crossover points in this set of plots as points where the supervised method (the green line) crosses with the CFT method (the blue line) in Figs. 4 and 3. This point shows the amount of data with clean labels needed for the supervised-only method to surpass the WS method that has access to the same labels as well as the LFs: the higher the crossover point, the more utility WS brings for that task.
We first conducted a crossover point analysis for the existing datasets from WRENCH [44] in Fig. 3, extending the results from [47]. We confirm that for most of these tasks, the crossover points are considerably smaller, less than 200 for 4 out of 6 tasks. Notably, these four datasets all have considerably smaller label cardinalities compared to the new datasets in BoxWRENCH (See Appendix D). These experiments verified the existing tasks are inadequate for benchmarking WS methods in realistic scnearios where WS holds significant utility over hand-labeling. We also performed the analysis on the named entity recognition datasets from WRENCH, with similar results (See Appendix H).
We then conducted the same experiments on our new datasets, and analyzed their crossover points in Fig. 4. For both Amazon31 and Banking77, the crossover points are beyond 1,000 clean labels. For Claude9, the validation set is smaller and even training on all available examples does not result in a crossover. Instead, the gap between the CFT and supervised-only methods remains 5% higher.
Figure 3: Crossover points on six existing WS benchmarks: AGNews, Semeval, Yelp, ChemProt, Trec, IMDB. The crossover points for these tasks are low, but for four out of the six tasks, the crossover points are less than 200, which is substantially lower than the crossover points in BoxWRENCH datasets.
Figure 4: Crossover points on three of our datasets: Amazon31, Banking77, and Claude9. For both Amazon31 and Banking77, the crossover points are beyond 1,000 clean labels. For Claude9, the validation set is smaller and even training on all available examples does not result in a crossover.
4.2 Improving LFs leads to higher crossovers
Previously, the LFs of ChemProt from Yu et al. [43], which are keywords-based, have a coverage of 0.86 on the test set, and a precision of 0.55 (i.e., the accuracy of on the set of examples where LM does not abstain).To improve the quality of the LF set, we made the following changes. First we changed keywords used in existing LFs. Next, we modified seven LFs by using the absolute difference in entity positions in the text as features. After making these changes, we had a coverage of 0.8102 and precision of 0.6321. When constructing these LFs, we consulted domain experts, (Jinoh Li, a Ph.D students in Neuroscience and Behavior at Harvard) that have sufficient Chemistry and Biology knowledge for insights on the LFs. The detailed modification is included in Fig. 13.
These modifications also resulted in higher accuracy for the continuously fine-tuned model given the same amount of labeled data. The accuracy crossover point is around 800 (in comparison to the old 600 as shown in Fig. 3), and the F1 score crossover point is around 1600, whereas the previous F1 score crossover points is around 800. ChemProt has high class imbalance, so the F1 score is an essential metric to consider.
This experiment demonstrates that more careful writing of LFs can yield higher crossover points. Current research benchmarks may suffer from suboptimal LF sets, and this example highlights how simple adjustments can significantly enhance results.
Figure 5: Crossover points on ChemProt. The new LF sets for ChemProt demonstrate a higher crossover point in both accuracy and F1-micro scores, reaching over 800 compared to the previous 600 as in Fig. 3. This shift suggests a notable improvement in performance.
4.3 Cross-lingual adaptability
For MASSIVE18, we consider the setting in which one has already trained a weakly supervised model based on the English version of the dataset and wishes to also be able to obtain a strong model for a different language. In doing so we consider the following approaches that leverage the portability of LFs which is a key advantage of WS deployments:
[…]
  • L2En-LFs, this method intends to provide a more realistic scenario for reusing the English LFs. In practical scenarios, we might have a set of English LFs $\text{LF}_{\text{EN}}$ and need to train a model in another language, L, without the resources to create additional LFs in a foreign language. In this case, we use DeepL [9] to first translate our non-English unlabeled data to English and then apply $\text{LF}_{\text{EN}}$ to obtain weak labels. Fig. 2 illustrates this pipeline.
  • L2En-Inference-Supervised. This is an important baseline where we evaluate whether the English model alone can be directly used on other languages by simply first translating test examples to English. In Fig. 6, the validation size for this method represents the amount of clean data used to train the English model.
In this section, we present a set of experiments to demonstrate another dimension of WS’s usefulness: adaptability. Specifically, we look at this through the ability to re-use existing LFs in a multilingual setting. Here we present the accuracy of $f_{\theta_{\text{WS}}}$, $f_{\theta_{\text{CFT}}}$, and $f_{\theta_{\text{SUP}}}$ with our methods in Fig. 6, where different suffixes are used with each method for clarity. We use L2En-Inference-Supervised as a baseline, representing the use of non-language-specific models. For any given task, this method translate the data back to English and then use the existing English model. Fig. 6 demonstrates that language-specific models consistently achieve higher accuracy.
Figure 6: Crossover points on the multilingual MASSIVE18 dataset. The comparison of the green solid line and yellow solid line showed the importance of having a language specific model. Surprisingly using our method from Fig. 2(blue solid line) is able to achieve the same or even better performance compare to the light blue dotted line which used the weak labels from MASSIVE18.
The more realistic setting L2En-LFs Fig. 2 using back-translations also achieve comparable accuracy with Oracle-LFs which rely on the bijective relation of the original MASSIVE dataset, implying the usefulness of applying WS in real-world multi-language application. In addition, among all the datasets, we did not create new LFs in another languages; instead, we leveraged the generalizability of the existing LFs. As a result, the higher crossover points benefits from WS here are almost for free. We further noted that without developing language-specific models, the translation costs would be prohibitively high, because translation costs would incur at each inference time. We also noted that WS boost the crossover points to over 1000 even with the “free” noisy weak labels from our method.
Finally, we introduce a realistic setup in which we take 20% of the unlabeled training data from Norwegian, translate it to English, and apply the LFs. We then attach the weak labels generate by MV to the untranslated version. Additionally, we take 80% of the unlabeled training data with weak labels from English and translate them into Norwegian. These steps, combined with a Norwegian BERT model, leads to $f_{\theta_{\text{WS}}}$. The results is presented in Fig. 9, where the crossover points is above 1500, showcasing just by using existing LFs can provide incredible performance.
For languages where foundation models are less powerful, the techniques used by L2En-LFs can help to re-use the LFs and save a great amount of clean labels. This is helpful, especially when rarer languages have a limited clean data or even unlabeled data.
4.4 Label Model
Table 2: Test accuracy (ChemPort, MASSIVE18)/F1 scores (Claude9) of supervised-only methods and CFT across different proportions of clean data used, LMs, and datasets. First and second best results are Bolded and underlined.
Claude9 ChemProt [^8] Here we use ChemProt with updated LFs. MASSIVE18
6.25% Validation Size
+Majority vote 0.153$\pm$0.026 0.625$\pm$0.018 0.811$\pm$0.004
+DawidSkene 0.153$\pm$0.025 0.610$\pm$0.020 0.797$\pm$0.010
+Snorkel 0.132$\pm$0.022 0.617$\pm$0.012 0.811$\pm$0.007
+FlyingSquid 0.109$\pm$0.005 0.606$\pm$0.021 0.751$\pm$0.011
+Supervised Only 0.109$\pm$0.005 0.504$\pm$0.030 0.753$\pm$0.014
12.5% Validation Size
+Majority vote 0.175$\pm$0.020 0.685$\pm$0.010 0.845$\pm$0.006
+DawidSkene 0.184$\pm$0.022 0.654$\pm$0.024 0.836$\pm$0.004
+Snorkel 0.157$\pm$0.016 0.649$\pm$0.015 0.845$\pm$0.005
+FlyingSquid 0.124$\pm$0.014 0.640$\pm$0.015 0.818$\pm$0.009
+Supervised Only 0.123$\pm$0.015 0.585$\pm$0.022 0.820$\pm$0.013
25% Validation Size
+Majority vote 0.345$\pm$0.114 0.726$\pm$0.021 0.863$\pm$0.004
+DawidSkene 0.362$\pm$0.107 0.712$\pm$0.014 0.863$\pm$0.007
+Snorkel 0.303$\pm$0.090 0.711$\pm$0.016 0.859$\pm$0.002
+FlyingSquid 0.170$\pm$0.025 0.696$\pm$0.010 0.856$\pm$0.009
+Supervised Only 0.165$\pm$0.052 0.684$\pm$0.023 0.855$\pm$0.010
50% Validation Size
+Majority vote 0.483$\pm$0.071 0.778$\pm$0.005 0.882$\pm$0.005
+DawidSkene 0.477$\pm$0.074 0.775$\pm$0.003 0.884$\pm$0.005
+Snorkel 0.462$\pm$0.065 0.767$\pm$0.009 0.885$\pm$0.005
+FlyingSquid 0.232$\pm$0.057 0.762$\pm$0.008 0.880$\pm$0.005
+Supervised Only 0.253$\pm$0.043 0.773$\pm$0.010 0.883$\pm$0.004
100% Validation Size
+Majority vote 0.582$\pm$0.036 0.820$\pm$0.004 0.899$\pm$0.002
+DawidSkene 0.572$\pm$0.038 0.820$\pm$0.005 0.893$\pm$0.002
+Snorkel 0.557$\pm$0.020 0.814$\pm$0.004 0.899$\pm$0.002
+FlyingSquid 0.352$\pm$0.037 0.813$\pm$0.006 0.894$\pm$0.005
+Supervised Only 0.347$\pm$0.020 0.816$\pm$0.007 0.898$\pm$0.003
In real-world applications, WS is typically integrated with different LMs to optimize performance. Several LMs are frequently employed in WS frameworks, including Majority Vote, Dawid-Skene [8], Snorkel [33], and FlyingSquid [13].
In our study, we conducted a comprehensive evaluation of WS using these LMs on our datasets. The performance of each LM was systematically assessed to determine its effectiveness in various scenarios. Detailed results of this evaluation are presented in Table 2, showcasing the comparative performance and highlighting the strengths of each model (See Appendix F for the full table with Amazon31 and Banking77).
For most of the new dataset, the CFT method outperforms the supervised-only method with a clear margin, especially in low-resource settings. Majority Vote and Dawid-Skene performed the best among the LM tested.
5 Limitations
There are several limitations to this work.
  • This work primarily focuses on text classification tasks, which are more common in practice; while other benchmarks such as WRENCH Zhang et al. [44] have both text classification and sequence-tagging tasks. Similar investigation and LF improvements could be confused on those tasks using our strategy and code base in further work.
  • This paper only primarily evaluated the performance of RoBERTa’s (some other BERT variant) EM. Similarly, this paper only tested the performance of some relatively simple LMs such as Majority Vote, Dawid-Skene, Snorkel, COSINE, and ARS2. More complex and recent models could be tested using the pipeline.
  • This paper fixed the number of steps and hyperparameter while training following settings from [47]. With proper early stopping and hyperparameter search, WS has the potential to achieve better results.
  • The MASSIVE18 dataset has one-to-one correspondence across languages, while in real-life scenarios, distribution shifts among different languages even on identical tasks. Thus multi-lingual dataset collected from real uses would be more effective in evaluating the performance
6 Conclusions
In this paper, we introduce BoxWRENCH, a benchmark that expands the evaluation of WS by addressing the limitations of existing benchmarks. By incorporating high-class cardinality, imbalance, and the need for domain expertise, BoxWRENCH  better reflects real-world data and tasks. Our results demonstrate that WS is highly effective, particularly in scenarios where traditional labeling is cost-prohibitive, and it achieves substantial performance gains before supervised methods surpass it at high label counts. We also establish that careful LF design and leveraging existing LFs in multilingual settings can significantly enhance WS applicability across diverse contexts. BoxWRENCH  sets a new standard for evaluating WS, with publicly released datasets, benchmarks, and tools to advance WS research and its practical deployment.
Acknowledgements.
We are grateful for the support of the NSF under CCF2106707 (Program Synthesis for Weak Supervision) and the Wisconsin Alumni Research Foundation (WARF). Neel Guha is supported by the Stanford Interdisciplinary Graduate Fellowship.
Appendix A Broader Social Impact of WS Benchmarking
Our benchmark aims to provide a platform to evaluate WS methods on more realistic datasets. Methods with successful performance are more likely to be useful in real applications, thus improving the effectiveness of WS and reducing the potential costs for practitioners wishing to train ML models. Of course, a potential negative societal impact is that if models are easier to train with WS, people with malicious intent can also train models at a relatively lower cost, leading to potentially harmful impacts.
Appendix B Dataset Licenses
Our license is CC BY 4.0 license and otherwise inherits the licensing of original datasets.
Appendix C LF Improvements for Chemprot
We described the details of how we improved the LFs for the Chemprot dataset in this document. Further details can be found in our codebase.
C.1 Original LFs
The original LFs have a coverage of 0.8637 and precision of 0.5512 on the covered data, with an accuracy of 0.4904 using Majority Vote and random tie-breaking (Wrench reported their accuracy in this way). These statistics are obtained from the test set.
We sampled a development set of size 250 from the training set, examined the definition of each label, and carefully reviewed examples of each label to understand the characteristics of the dataset.
Example LFs are shown below, with the full set found at lable_function/chemprot.
# chemprot functions examples:
#0
@labeling_function()
def lf_amino_acid(x):
    return 0 if ’amino acid’ in x.text.lower() else ABSTAIN
#19
## Cofactor
@labeling_function()
def lf_cofactor(x):
    return 7 if ’cofactor’ in x.text.lower() else ABSTAIN
C.2 LF Improvement Details
We first started by adding space around or before the keywords in some LFs: “activat”, “increas”, “reduc”, “antagon”, “transport”, “catalyz”, “produc”, and “not”. This is because for keywords such as “not”, they might be triggered by words like “notable”.
We also removed LFs with low accuracy on the development set. For example, we removed the function lf_induce, as the word "induce" is too general.
Additionally, we developed a utility function, chemprot_enhanced, to extend the chemprot dataframe with two more columns: entity1_index and entity2_index. We improved our LFs to utilize these indices to check whether certain words occur between or near the two entities.
After these improvements, on the development set, our coverage dropped to 0.828, but the accuracy for covered data increased to 0.5942. Accuracy with Majority Vote and random tie-breaking rose to 0.508.
Figure 7: Comparison of Chemprot results
Appendix D Cardinality for WS tasks
Table 3: Number of classes for datasets in WRENCH (top) compared to the new tasks in BoxWRENCH(bottom)
Dataset Number of Classes
IMDB 2
ChemProt 10
TREC 6
Yelp 2
SemEval 9
AGNews 4
Banking77 77
Claude9 9
MASSIVE18 18
MASSIVE60 60
Amazon31 31
Appendix E Reproducing results from Zhu et al. [47].
Here, we show how using our codebase, we can successfully reproduce the results from Zhu et al. [47].
Dataset N Implementation Supervised Weakly Supervised CFT
AGNEWS 50 Zhu et al. [47] 0.880 0.872 0.882
5 Zhu et al. [47] 0.770 0.840 0.841
50 Ours 0.875 $\pm$ 0.007 0.859 $\pm$ 0.010 0.871 $\pm$ 0.017
5 Ours 0.769 $\pm$ 0.040 0.863 $\pm$ 0.009 0.820 $\pm$ 0.030
Yelp 50 Zhu et al. [47] 0.950 0.820 0.910
5 Zhu et al. [47] 0.740 0.760 0.840
50 Ours 0.947 $\pm$ 0.006 0.868 $\pm$ 0.043 0.943 $\pm$ 0.075
5 Ours 0.767 $\pm$ 0.056 0.843 $\pm$ 0.072 0.915 $\pm$ 0.014
IMDb 50 Zhu et al. [47] 0.880 0.818 0.864
5 Zhu et al. [47] 0.705 0.795 0.797
50 Ours 0.868 $\pm$ 0.030 0.846 $\pm$ 0.023 0.889 $\pm$ 0.007
5 Ours 0.630 $\pm$ 0.064 0.819 $\pm$ 0.027 0.794 $\pm$ 0.054
TREC 50 Zhu et al. [47] 0.930 0.680 0.940
5 Zhu et al. [47] 0.630 0.640 0.840
50 Ours 0.911 $\pm$ 0.014 0.678 $\pm$ 0.097 0.910 $\pm$ 0.013
5 Ours 0.603 $\pm$ 0.046 0.662 $\pm$ 0.036 0.815 $\pm$ 0.040
ChemProt 50 Zhu et al. [47] 0.720 0.550 0.730
5 Zhu et al. [47] 0.420 0.510 0.590
50 Ours 0.707 $\pm$ 0.0163 0.583 $\pm$ 0.012 0.737 $\pm$ 0.0069
5 Ours 0.420 $\pm$ 0.023 0.518 $\pm$ 0.030 0.573 $\pm$ 0.027
SemEval 50 Zhu et al. [47] 0.862 0.820 0.910
5 Zhu et al. [47] 0.720 0.760 0.840
50 Ours 0.855 $\pm$ 0.0037 0.837 $\pm$ 0.016 0.916 $\pm$ 0.077
5 Ours 0.747 $\pm$ 0.021 0.836 $\pm$ 0.006 0.868 $\pm$ 0.006
Appendix F Extended Table for Section 4.4
Table 4 extends the results in Section 4.4, showing additional results for Amazon31 and Banking77.
Table 4: Additional test accuracy for Amazon31 and Banking77. Best results are Bolded.
6.25% Validation 12.5% Validation 25% Validation 50% Validation 100% Validation
Banking77
+Majority vote 0.5652$\pm$0.0262 0.6665$\pm$0.0152 0.7519$\pm$0.0150 0.8218$\pm$0.0098 0.8654$\pm$0.0023
+Supervised Only 0.2362$\pm$0.0155 0.3777$\pm$0.0199 0.5705$\pm$0.0176 0.7522$\pm$0.0041 0.8488$\pm$0.0065
Amazon31
+Majority vote 0.6850$\pm$0.0046 0.7103$\pm$0.0066 0.7429$\pm$0.0028 0.7706$\pm$0.0032 0.7923$\pm$0.0016
+Supervised Only 0.6603$\pm$0.0078 0.6999$\pm$0.0056 0.7357$\pm$0.0026 0.7673$\pm$0.0033 0.7902$\pm$0.0012
Appendix G MASSIVE60 and Using English for Augmentation
This section contains the results for MASSIVE60 (Fig. 9), as well as an exploratory approach for leveraging non-English data as additional weak supervision for improving performance on MASSIVE-En. Section 4.3 (Fig. 9).
Figure 8: MASSIVE60
Figure 9: Using English text as an augmented data for Norwegian BERT
Appendix H NER datasets
We included additional experiments related to named entity recognition datasets (BC5CDR[24], NCBI-Disease[11], CoNLL-03[36], OntoNotes 5.0[42]) at Fig. 11. Overall, we obeserve that crossover points are relatively small but note that each sequence has multiple label entities, which would further increase the cost of manual labeling compared to sequence classification tasks. Finding tasks and LFs for NER that have higher crossover points would be an interesting avenue for future work.
Figure 10: Crossoverpoints for additional NER task.
Figure 11: Crossoverpoints for additional NER task.
Appendix I Cosine + ARS2
We experimented with an additional weak supervision baseline with the same experimental pipeline. We include results for COSINE and ARS2 in Fig. 12. We used both of the learning methods with a RoBERTa/Legal BERT backbone on the WS-only pipeline. The other setups were kept the same. The ARS2 results follow a similar trend to our previous results, with almost identical cross-over points. COSINE performance is slightly worse on Amazon31 and Banking77, and is slightly better on Claude9, when compared with the supervised-only method.
Figure 12: Crossover points using ARS2 and Cosine as end models.
Appendix J F1 scores for all the dataset
Since some of the datasets are highly imbalanced, we provide other metrics (micro, macro, weighted F1 scores) for those Table 5, providing a more complete picture of performance across imbalanced classes. The inclusion of these metrics does not change any of our conclusions.
Table 5: Accuracy and F1 scores for different datasets and methods with varying validation sizes (VS). Best results are bolded.
6.25% VS 12.5% VS 25% VS 50% VS 100% VS
Amazon31
    +CFT (Accuracy) 0.685 0.711 0.742 0.769 0.791
    +Supervised Only (Accuracy) 0.660 0.701 0.736 0.766 0.790
    +CFT (F1_micro) 0.685 0.711 0.742 0.769 0.791
    +Supervised Only (F1_micro) 0.660 0.701 0.736 0.766 0.790
    +CFT (F1_macro) 0.683 0.710 0.741 0.768 0.791
    +Supervised Only (F1_macro) 0.659 0.702 0.736 0.766 0.790
    +CFT (F1_weighted) 0.682 0.709 0.740 0.767 0.790
    +Supervised Only (F1_weighted) 0.658 0.701 0.735 0.765 0.789
Banking77
    +CFT (Accuracy) 0.551 0.676 0.751 0.820 0.866
    +Supervised Only (Accuracy) 0.234 0.378 0.567 0.753 0.846
    +CFT (F1_micro) 0.551 0.676 0.751 0.820 0.866
    +Supervised Only (F1_micro) 0.234 0.378 0.567 0.753 0.846
    +CFT (F1_macro) 0.497 0.650 0.742 0.818 0.865
    +Supervised Only (F1_macro) 0.167 0.318 0.530 0.745 0.844
    +CFT (F1_weighted) 0.497 0.650 0.742 0.818 0.865
    +Supervised Only (F1_weighted) 0.167 0.318 0.530 0.745 0.844
Claude9
    +CFT (Accuracy) 0.904 0.906 0.906 0.914 0.914
    +Supervised Only (Accuracy) 0.907 0.908 0.906 0.911 0.913
    +CFT (F1_micro) 0.904 0.906 0.906 0.914 0.914
    +Supervised Only (F1_micro) 0.907 0.908 0.906 0.911 0.913
    +CFT (F1_macro) 0.149 0.177 0.351 0.493 0.558
    +Supervised Only (F1_macro) 0.110 0.120 0.173 0.256 0.349
    +CFT (F1_weighted) 0.874 0.877 0.895 0.910 0.916
    +Supervised Only (F1_weighted) 0.864 0.866 0.870 0.884 0.900
ChemProt
    +CFT (Accuracy) 0.637 0.681 0.724 0.776 0.819
    +Supervised Only (Accuracy) 0.595 0.606 0.605 0.614 0.613
    +CFT (F1_micro) 0.637 0.681 0.724 0.776 0.819
    +Supervised Only (F1_micro) 0.595 0.606 0.605 0.614 0.613
    +CFT (F1_macro) 0.453 0.523 0.572 0.640 0.710
    +Supervised Only (F1_macro) 0.405 0.410 0.412 0.409 0.412
    +CFT (F1_weighted) 0.623 0.669 0.715 0.770 0.816
    +Supervised Only (F1_weighted) 0.573 0.581 0.581 0.587 0.586
Appendix K SciBERT
We used in-domain SciBERT[4], a pretrained language model for scientific text to test whether the crossover point is still consistent. The crossover point is even higher in this case, suggesting the usefulness of weak supervision, see Fig. 13.
Figure 13: SciBERT Results
Appendix L Comparison with or using LLMs
LLMs have demonstrated strong zero-shot or few-shot capabilities, one may also curious about how SOTA LLMs perform on our datasets. We sampled 250 data points from our test set, utilized API calls with GPT-4o-2024-08-06 for each example, and recorded the accuracy. We attached the cross-over points graph with this baseline (See Fig. 14). The LLM baseline performed poorly in tasks requiring domain-specific knowledge, such as Claude9 and ChemProt. It performed well on Amazon31, possibly because the test set may be included in the public amazon review datasets, which GPT may have been trained on.
Figure 14: Using LLM as baseline
We also prompted GPT-4o as if it was a domain expert (e.g. “you are an expert in legal document classification and label function writing”) and asked it to create keywords-based LFs based upon a development set. The full list of our prompts is included in our codebase. While the coverage of the GPT-4o-generated label functions is higher, the precision of the LFs is generally lower (see Table 6). However, we note that the generated labeling functions (LFs) are acceptable, especially considering the reduction in time cost and the potential for full automation. Future work could involve analyzing a multi-agent scenario that automates and reinforces the LF generation process.
Table 6: Coverage and Precision Comparison Between Our LFs and LLM-Generated LFs
Dataset Our Coverage Our Precision LLM Coverage LLM Precision
Amazon31 0.63 0.70 0.99 0.26
Banking77 0.58 0.81 0.98 0.62
ChemProt 0.82 0.70 0.96 0.56
Claude9 1.00 0.91 0.91 0.15
Appendix M Miscellaneous
M.1 Links to datasets & metadata
We noted that all the datasets that we used are based on previously publicly published datasets. The license and links are mentioned in Appendix B. In addition to the original link for the datasets mentioned in Appendix B. We also provide our own usage of the datasets on GoogleDrive.
M.2 Dataset Format
The dataset uses the same format as the WRENCH [44] benchmark. For each dataset directory, there are four .json files for training data, validation data, test data, and labels respectively. For the label file, the labels are organized in the following format:
    Label index: Label name,
For the dataset files, the data points are organized in the following format:
    Data index: {
        "labels": Label index,
        "weak_labels": [-1, -1, ...],
        "data": {
            "text": Data Content,
            More content depending on the datasets
M.3 Structured Metadata
We also provide our datasets on HuggingFace, and the metadata are contained in the README.md for each dataset.